OpenStudy (anonymous):

The strength of an electric field at 5.0 cm from a point charge is 100.0 N/C. What is the magnitude of the source charge?

OpenStudy (michele_laino):

here we can apply this formula: \[\Large E = K\frac{Q}{{{d^2}}}\] where \(E=100\), \(K=9 \cdot 10^9\), and \(d=0.05\) meters

OpenStudy (anonymous):

So the equation would be \[100=9\times10^9 \frac{ Q }{ (0.05^2) }\] ?

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

here is the next step: \[\huge Q = \frac{{E{d^2}}}{K} = \frac{{{{10}^2} \cdot 25 \cdot {{10}^{ - 4}}}}{{9 \cdot {{10}^9}}} = ...?\]

OpenStudy (anonymous):

So then \[\frac{ .25 }{ 9\times10^9 }=2.7^-11\]?

OpenStudy (anonymous):

This is the result that I got in the end: \[Q=2.7\times10^-11C\] Is that correct?

OpenStudy (michele_laino):

yes! It is correct!

OpenStudy (anonymous):

Thank you! (: